How do you solve this problem? not sure where to start
9 Tan^2 x + 5 = 8

Answer by lwsshak3(11628)

9 Tan^2 x + 5 = 8
9 Tan^2 x=3
Tan^2 x=3/9=1/3
Take sqrt of both sides
Tan x=1/√3
x=30º and 210º (in quadrants I and III where tan>0) (0 < x < 360º)